OpenStudy (anonymous):

integral 3x + 1 squared dx

OpenStudy (cwrw238):

you can do this by expanding and integrating term by term (3 x + 1)^2 = 9x^2 + 6x + 1 then use the standard algebraic tntegratio INT a x^n dx = a x^(n+1) ------- n + 1 and then add the constant of integration C

OpenStudy (anonymous):

You can also use a substitution approach, letting \(u=3x+1\), then \(\dfrac{1}{3}du=dx\), so \[\int(3x+1)^2~dx=\frac{1}{3}\int u^2~du\]